Tartan Industries currently has total capital equal to $4 million, has zero debt, is in the 40% federal-plus-state tax bracket, has a net income of $4 million, and distributes 40% of its earnings as dividends. Net income is expected to grow at a constant rate of 6% per year, 210,000 shares of stock are outstanding, and the current WACC is 14.00%.

The company is considering a recapitalization where it will issue $5 million in debt and use the proceeds to repurchase stock. Investment bankers have estimated that if the company goes through with the recapitalization, its before-tax cost of debt will be 11% and its cost of equity will rise to 14.5%.

a. What is the stock's current price per share (before the recapitalization)? Round your answer to the nearest cent. Do not round intermediate steps.

b. Assuming that the company maintains the same payout ratio, what will be its stock price following the recapitalization? Assume that shares are repurchased at the price calculated in part a. Round your answer to the nearest cent. Do not round intermediate steps.
Answer #1

a.Dividend = net income*payout ratio/shares outstanding = 4000000*0.4/210000=7.619

As per DDM
Price = recent dividend* (1 + growth rate )/(cost of equity - growth rate)
Price = 7.619 * (1+0.06) / (0.14 - 0.06)
Price = 100.95

b

New net income = old net income-debt*interest rate *(1-tax rate)

=4000000-5000000*0.11*(1-0.4)=3670000

New shares = old shares-debt/price

=210000-5000000/100.95 =160470.53

New dividend = New net income*payout ratio/new shares = 3670000*0.4/160470.53=9.148

As per DDM
Price = recent dividend* (1 + growth rate )/(cost of equity - growth rate)
Price = 9.148 * (1+0.06) / (0.145 - 0.06)
Price = 114.08